Circular No. 5771

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1993K IN NGC 2223 M. Hamuy, Cerro Tololo Interamerican Observatory, obtained a low-

dispersion spectrogram (range 320-750 nm) of SN 1993K with the CTIO 4.0-m telescope on Apr. 21.05 UT. These data show that strong, P-Cyg emission lines of H I and Fe II have developed, confirming earlier suspicions that this object is a type-II event (cf. IAUC 5733, 5736). The spectrum closely resembles that of the type-IIP supernova 1990E obtained 17 days after discovery (see Schmidt et al. 1993, A.J., in press).

SUPERNOVA 1993J IN NGC 3031 A. Tomaney and A. Crotts, Columbia University, communicate: “Ten-

hr pre-explosion CCD exposures in both B and Hα (7.5-nm bandpass) of the field of SN 1993J were obtained using the McDonald Observatory 2.1m telescope during 1990 Feb. and Mar. The source is within 0′′.4 of the SN 1993J position given by Richmond on IAUC 5737. Images with seeing 2′′.3 FWHM or better show the object to be extended (FWHM of 3′′.0 and 3′′.6 in B and Hα, respectively). Lucy-deconvolution of the images yields FWHM of 2′′.9 in Hα and 2′′.6 in B for the source, compared to 1′′.3 for nearby stars, but reveals no substructure within the source. The Hα flux in a 4′′-radius aperture was 2.92 (± 0.05) × 10^{−17} erg s^{−1} cm^{−2} A^{−1}, in good agreement with the 1992 Jan. measurement by Magnier et al. (IAUC 5741). This is roughly equivalent to an R magnitude of 19.3, very close to the Rband measure in 1990 Mar. by Perelmuter (IAUC 5736), indicating that no strong Hα emission was associated with the source at that time. The source sits on a ridge of diffuse, unresolved starlight ∼ 3′′ wide and 20′′ long with p.a. 40^{o}, appearing stronger in B than Hα. Differential photometry using nearby stars yields a source magnitude of B = 20.44 ± 0.03 in a 4′′-radius aperture, versus B = 21.2 in a 1′′ aperture, the latter being closer to previous estimates. CCD saturation of stars brighter than mag 16 necessitated a calibration of fainter stars in the field. The magnitudes quoted here are made relative to a star 23′′.2 west and 56′′.5 south of the supernova, where we measure 2.66 (± 0.01) × 10^{−16} erg s^{−1} cm^{−2} A^{−1} for the Hα flux and B = 18.95 ± 0.03.”

1993 April 21 Daniel W. E. Green